# Enhanced pedestal transport driven by edge collisionality on Alcator C-Mod and its role in regulating H-mode pedestal gradients

### v1

@@ 6 Discussion
As mentioned in Section 3, pedestal stiffness and critical gradient-like behavior in the edge of Alcator C-Mod is not a new finding. Given C-Mod’s proximity in parameter space to next-generation devices [56, 57, 58], it is imperative to develop improved understanding for the conditions under which this phenomenon occurs. This paper proposes collisional-driven transport as the mechanism by which a pedestal at its natural density responds to increased neutral flux and regulates its density gradient. As $S_{\mathrm{ion}}^{\mathrm{sep}}$ increases, so does $n_{e}^{\mathrm{sep}}$. Higher $S_{\mathrm{ion}}^{\mathrm{sep}}$ corresponds to an increase in $n_{0}^{\mathrm{sep}}$ as well, which presumably increases plasma thermal sinks, dropping $T_{e}^{\mathrm{sep}}$. The concomitant drop in $T_{e}^{\mathrm{sep}}$ ultimately drives pedestals that are dense with neutrals to higher collisionalities. As the collisionality at the separatrix transitions past a critical value, so does $\alpha_{t}$, and the transport changes in nature. At a critical gradient, increased fueling increases transport. It is this additional resistive transport mechanism that locks the gradient and drives the system non-diffusively. Given that these plasmas are all in the EDA regime, it is natural to question whether pedestal regulation via separatrix collisionality is inherently a feature of this regime. It has been shown on Alcator C-Mod that in EDA H-modes, the density fluctuations characteristic of the quasi-coherent mode (QCM) may be responsible for increased particle transport [59]. Analysis to evaluate possible links between the QCM and the type of transport described here are ongoing.
@@
It is important to note that this work says little about how a pedestal might reach a critical gradient to begin with and what types of devices might be susceptible to reaching one. Two characteristics of Alcator C-Mod, both hinted at above, appear important. The first of these is a neutral effect, which affects particle transport via indirect modification of collisionality. Having operated at the highest of plasma densities, C-Mod is proposed to be the device closest to the opaqueness predicted on both SPARC and ITER [15]. Inability to fuel in the pedestal will push the ionization front radially outward, increasing collisionality at the separatrix or in the near SOL, thereby exacerbating the gradient regulation via resistive transport. This may very well occur at high densities in future machines.
@@
The second effect is related to the poloidal distribution of neutrals on C-Mod. It existed in the so-called main chamber (MC) recycling regime [23], consistent with large radial particle fluxes observed at the OMP. Measurements from a gas puff imaging (GPI) system as well as from a stochastic model imply that this is linked to large filaments called “blobs” propagating radially outward, depositing a large number of particles and heat to the main chamber walls [60]. Little is known about the formation of these blobs, but it is likely that this is strongly linked to conditions at the pedestal. In either case, large particle fluxes to the MC walls means high main chamber neutral density, setting up a plasma environment conducive to the rise of RBMs at or near the separatrix. While it is unclear if the cause of high MC fueling on C-Mod is its high density or rather the close-fitting wall, these features will also be present on next-generation devices, and MC recycling may continue to dominate fueling.
@@
That these C-Mod pedestals exhibit what appears to be critical gradient behavior has a number of implications for new fusion devices. To realize sufficient fusion gain, reactors will require high core densities. These must also be compatible with robust pedestals for good confinement. This work shows that good confinement can be lost when $P_{\mathrm{net}}$ drops below a critical value, $P_{\mathrm{net}}^{\mathrm{crit}}$, even while remaining in H-mode. For the set of discharges analyzed, this critical value corresponds almost exactly to $P_{\mathrm{th}}^{\mathrm{L-H}}$, calculated from the Martin scaling [61]. Since next-generation devices like SPARC and ITER, and later reactors, require high density and large plasma volumes, it is possible that they too will operate at $P_{\mathrm{net}}$ marginal to their corresponding $P_{\mathrm{th}}^{\mathrm{L-H}}$ [62, 63, 64]. It remains to be seen whether the observed transition to a collisional transport regime is inherently linked to the power threshold in a quantitative way in general, or just by coincidence on C-Mod in this parameter range. Once again, we note that this paper includes analysis only of H-modes that are not actively gas puffed. While intrinsic wall-recycling may sound attractive, it means density control is substantially more challenging, or at least that there is a strong (although perhaps indirect) link between $P_{\mathrm{net}}$ and plasma fueling.
@@ 7 Conclusions and future work
This work presents new analysis of archival data from Alcator C-Mod. It deploys the routinely used ETS together with the less-frequently analyzed LYMID data to study the role of fueling and transport on pedestals in a high density, opaque edge plasma. Using inferences of ionization rates, it is observed that $n_{e}^{\mathrm{sep}}$ is a strong function of $S_{\mathrm{ion}}$, continuing to increase even at high $S_{\mathrm{ion}}$. $n_{e}^{\mathrm{ped}}$, on the other hand, stagnates, implying a saturated and even shrinking $\nabla n_{e}$. Comparing this gradient with $\Gamma_{D}$ shows a non-linear relationship, only explainable with varying $D_{\mathrm{eff}}$, despite relatively fixed shape and $q_{95}$. Inspection of profiles inside the pedestal top shows that the phase space at this radial location is best categorized by $p_{e}$ and $\nu^{*}$. Highly collisional pedestals are also found to have lower values of $\alpha_{\mathrm{MHD}}$, consistent with worsening H-mode quality and thus, $H_{98}$. It is found that $\nu^{*}_{\mathrm{sep}}$ is most influential in determining the large rise in $D_{\mathrm{eff}}$ across the pedestal and the subsequent flattening of the edge $n_{e}$ profile. Initial analysis links this to changes in $\alpha_{t}$, which mediates the transition betwen DW and RBM turbulence.
@@
We supplement these experimental findings of changes to $D_{\mathrm{eff}}$ through modeling of selected discharges with the widely used edge code, SOLPS-ITER. High fidelity calculation of particle and heat sources, in addition to experimental 1D constraints, lends confidence to conclusions about how particle and heat transport must change to build up pedestal-gradients. It highlights the importance of self-consistent modeling of both sources and transport in the development of models for the pedestal. In particular, it allows for inference of $\chi_{e}$ (and $\chi_{i}$ to the degree that we trust the $\tau_{i}=1$ assumption). Large growth in $\chi_{e}$ bolsters the claim that changes to the fundamental transport properties in these pedestals may be at play. Though not highlighted here, work has begun to understand how the poloidal distribution of neutrals may validate or limit the 1D inferences of $D_{\mathrm{eff}}$ from experiment. This, as well as understanding the contributions from different atomic processes to the neutral density in the pedestal will be important steps in developing a model for the interplay between fueling and transport in the pedestal.
@@
It has become evident that reactors will not be able to tolerate Type-I ELMs [64, 65, 66]. The non-ELMing behavior of the EDA H-mode, and other similarly non-ELMing modes, highlights the importance in understanding pedestal structure, and the mechanisms that determine it. While it is unclear if the EDA regime is achievable on reactors given that they will operate at high $T_{e}^{\mathrm{ped}}$ and thus, low $\nu^{*}_{\mathrm{ped}}$, some ELM-free solution is required. The analysis presented here works to improve understanding of how a pedestal may reach a transport limit before it is limited by less benign MHD events. It furthermore suggests that transport throughout the pedestal is more closely linked to the collisionality at the separatrix rather than at the pedestal top. Even if $\nu^{*}_{\mathrm{ped}}$ were to remain low, as long as $\nu^{*}_{\mathrm{sep}}$ remains high enough, transport regulation of pedestal gradients may persist, ensuring ELM-free operation. Given that reactors may operate at high $n_{e}^{\mathrm{sep}}$ for divertor survivability, the requisite $\nu^{*}_{\mathrm{sep}}$ may not altogether be out of reach. High edge collisionality may also help with turbulence spreading, potentially widening near-SOL widths and working towards an integrated core-edge exhaust solution.

### v2

@@ 6 Discussion
As mentioned in Section 3, pedestal stiffness and critical gradient-like behavior in the edge of Alcator C-Mod is not a new finding. Given C-Mod’s proximity in parameter space to next-generation devices [64, 65, 66], it is imperative to develop improved understanding for the conditions under which this phenomenon occurs. This paper proposes collisionally-driven transport as the mechanism by which a pedestal at its natural density responds to increased neutral flux and regulates its density gradient. As shown in Figure 4, as $S_{\mathrm{ion}}^{\mathrm{sep}}$ increases, so does $n_{e}^{\mathrm{sep}}$. Higher $S_{\mathrm{ion}}^{\mathrm{sep}}$ corresponds to an increase in $n_{0}^{\mathrm{sep}}$ as well, which presumably increases plasma thermal sinks, dropping $T_{e}^{\mathrm{sep}}$. The concomitant drop in $T_{e}^{\mathrm{sep}}$ ultimately drives pedestals that are dense with neutrals to higher collisionalities. Figures 8 and 9 show that as the collisionality at the separatrix increases, so does $\alpha_{t}$, as well as $D_{\mathrm{eff}}$, likely as a result of a change in the nature of the transport. Figure 5 shows that when the gradient reaches a certain value, increased fueling increases transport, and at large enough values, the gradient begins to shrink. This additional resistive transport mechanism necessitates the use of varying diffusive transport coefficients. Given that these plasmas are all in the EDA regime, it is natural to question whether pedestal regulation via separatrix collisionality is inherently a feature of this regime. It has been shown on Alcator C-Mod that in EDA H-modes, the density fluctuations characteristic of the quasi-coherent mode (QCM) may be linked to increased particle transport [67]. Analysis to evaluate possible links between the QCM and the type of transport described here are ongoing.
@@
It is important to note that this work says little about how a pedestal might reach a critical gradient to begin with and what types of devices might be susceptible to reaching one. Two characteristics of Alcator C-Mod, both hinted at above, appear important. The first of these is a neutral effect, which affects particle transport via indirect modification of collisionality. Having operated at the highest of plasma densities, C-Mod is proposed to be the device closest to the opaqueness predicted on both SPARC and ITER [15]. Inability to fuel inside the pedestal will push the ionization front radially outward, increasing collisionality at the separatrix or in the near SOL, thereby exacerbating the gradient regulation via resistive transport. This may very well occur at high densities in future machines. It is important to note that in terms of typical length scales, e.g. the poloidal gyroradius, the pedestal width, or the machine size, Alcator C-Mod is in a somewhat different parameter regime than future reactors. When considering dimensionless quantities, however, like opaqueness, or the gyroradius normalized to the minor radius, $\rho^{*}$, linked to both neoclassical and turbulent transport, this is not entirely the case. As for $\nu^{*}$, the parameter identified in this work as responsible for enhanced pedestal transport, large radial variation in this value makes it difficult to make conclusions about reactor-relevance, as noted in Section 4.2.
@@
The second effect is related to the poloidal distribution of neutrals on C-Mod. It existed in the so-called main chamber recycling regime [23], consistent with large radial particle fluxes observed at the OMP. Measurements from a gas puff imaging system as well as from a stochastic model imply that this is linked to large filaments called “blobs" propagating radially outward, depositing a large number of particles and heat to the main chamber walls [68]. In either case, large particle fluxes to the MC walls means high main chamber neutral density, setting up a plasma environment conducive to the rise of RBMs at or near the separatrix. While it is unclear if the cause of high MC fueling on C-Mod is its high density or rather the close-fitting wall, these features will also be present on next-generation devices, and MC recycling may continue to comprise a large fraction of the fueling dynamics.
@@
That these C-Mod pedestals exhibit what appears to be a limit to gradient growth has a number of implications for new fusion devices. To realize sufficient fusion gain, reactors will require high core densities. These must also be compatible with robust pedestals for good confinement. This work shows that good confinement can be lost when $P_{\mathrm{net}}$ drops below a critical value, $P_{\mathrm{net}}^{\mathrm{crit}}$, even while remaining in H-mode. For the set of discharges analyzed, this critical value corresponds almost exactly to L-H power threshold, $P_{\mathrm{th}}^{\mathrm{L-H}}$, calculated from the Martin scaling [69]. This scaling has shown good agreement with power threshold experiments on C-Mod [70]. Across a wide range in $I_{p}$ and $\overline{n}_{e}$, some deviations from the scaling exist at low and high $I_{p}$ and $\overline{n}_{e}$, but for $I_{p}$ = 0.8 MA, the data matches the scaling [71]. Since next-generation devices like SPARC and ITER, and later reactors, require high density and large plasma volumes, it is possible that they too will operate at $P_{\mathrm{net}}$ marginal to their corresponding $P_{\mathrm{th}}^{\mathrm{L-H}}$ [70, 71, 33].
It remains to be seen whether the observed transition to a collisional transport regime dictated by a collisional separatrix is inherently linked to $P_{\mathrm{th}}^{\mathrm{L-H}}$ in a quantitative way in general, or just by coincidence on C-Mod in this parameter range. Reactors, however, will likely require such a collisional separatrix with view of divertor handling [72], even if the pedestal is collisionless. Based on projections for $n_{e}^{\mathrm{sep}}$ and $T_{e}^{\mathrm{sep}}$ found in [73, 74], $\nu^{*}_{\mathrm{sep}}$ on SPARC may range anywhere from below 1 and up to 10. Operation at even higher $n_{e}^{\mathrm{sep}}$ might bring this value closer to values observed here. Additionally, the discharges analyzed here do not include seeded impurities. The lack of impurities allows for an easier estimate of $Z_{\mathrm{eff}}$, which figures into $\nu^{*}$, facilitating the analysis in this work. This may differ, however, from a reactor scenario where impurities may be used to access divertor detachment. We also note that this paper only includes analysis of H-modes that are not actively gas puffed, even with D gas. While intrinsic wall-recycling may sound attractive, it means density control is substantially more challenging, or at least that there is a strong (although perhaps indirect) link between $P_{\mathrm{net}}$ and plasma fueling.
@@ 7 Conclusions and future work
This work presents new analysis of archival data from Alcator C-Mod. It deploys the routinely used ETS together with the less-frequently analyzed LYMID data to study the role of fueling and transport on pedestals in a high density, opaque plasma edge. Using inferences of ionization rates, it is observed that $n_{e}^{\mathrm{sep}}$ is a strong function of $S_{\mathrm{ion}}$, continuing to increase even at high $S_{\mathrm{ion}}$. $n_{e}^{\mathrm{ped}}$, on the other hand, stagnates, implying a saturated and even shrinking $\nabla n_{e}$. Comparing this gradient with $\Gamma_{D}$ shows a non-linear relationship, only explainable with varying $D_{\mathrm{eff}}$, despite relatively fixed shape and $q_{95}$. Inspection of profiles inside the pedestal top shows that the phase space at this radial location is best categorized by $p_{e}$ and $\nu^{*}$. Highly collisional pedestals are also found to have lower values of $\alpha_{\mathrm{MHD}}$, consistent with worsening H-mode quality and thus, $H_{98}$. It is found that $\nu^{*}_{\mathrm{sep}}$ is most influential in determining the large rise in $D_{\mathrm{eff}}$ across the pedestal and the subsequent flattening of the edge $n_{e}$ profile. Initial analysis links this to changes in $\alpha_{t}$, which mediates the transition betwen DW and RBM-dominated turbulence.
@@
We supplement these experimental findings of changes to $D_{\mathrm{eff}}$ through modeling of selected discharges with the widely used edge code, SOLPS-ITER. High fidelity calculation of particle and heat sources, in addition to experimental 1D constraints, lends confidence to conclusions about how particle and heat transport must change to build up pedestal-gradients. Given the coupled nature of sources and transport demonstrated in this work, it recommends the development of models for the pedestal to include self-consistent accounting of both sources and transport. In particular, it allows for inference of $\chi_{e}$ (and $\chi_{i}$ to the degree that we trust the $\tau_{i}=1$ assumption). Large growth in $\chi_{e}$ bolsters the claim that changes to the fundamental transport properties in these pedestals may be at play. Though not highlighted here, work has begun to understand how the poloidal distribution of neutrals may validate or limit the 1D inferences of $D_{\mathrm{eff}}$ from experiment. This, as well as understanding the contributions from different atomic processes to the neutral density in the pedestal, will be important steps in developing a model for the interplay between fueling and transport in the pedestal.
@@
It has become evident that reactors will not be able to tolerate Type-I ELMs [33, 75, 76]. The non-ELMing behavior of the EDA H-mode, and other similarly non-ELMing modes, highlights the importance in understanding pedestal structure, and the mechanisms that determine it. While it is unclear if the EDA regime is achievable on reactors given that they will operate at high $T_{e}^{\mathrm{ped}}$ and thus, low $\nu^{*}_{\mathrm{ped}}$, some ELM-free solution is required. The analysis presented provides a useful experimental basis for how a pedestal may reach a transport limit before it is limited by less benign MHD events. It furthermore suggests that transport in the pedestal can be closely linked to the collisionality at the separatrix, even more so than at the pedestal top. Even if $\nu^{*}_{\mathrm{ped}}$ were to remain low, as long as $\nu^{*}_{\mathrm{sep}}$ remains high enough, transport regulation of pedestal gradients may persist, ensuring ELM-free operation. Given that reactors may operate at high $n_{e}^{\mathrm{sep}}$ for divertor survivability, the requisite $\nu^{*}_{\mathrm{sep}}$ may not altogether be out of reach. As has been observed previously, high edge collisionality may also help widen near-SOL widths [77, 78], helping work towards an integrated core-edge exhaust solution.